Consider a corporation whose taxable income without state income tax is

                                                                                                                                                                                                                                                                                             

Gross revenue

​$3,700,000

All expenses

​$1,600,000

If the marginal federal tax rate is​ 35% and the marginal state rate is​ 5%, compute the combined state and federal taxes using the two methods.

State taxable income is ​$_________. ​(Round to the nearest​ dollar.)

State income tax is ​$_________. ​(Round to the nearest​ dollar.)

Federal taxable income is ​$____________. ​(Round to the nearest​ dollar.)

Federal income tax is ​$__________. ​(Round to the nearest​ dollar.)

Combined state and federal tax is ​$_________. ​(Round to the nearest​ dollar.)

Combined tax rate is _____​%. ​(Round to two decimal​ places.)

Combined state and federal tax using combined tax rate is ​$___________. ​(Round to the nearest​ dollar.)
Answer #1

1. State taxable income

= 3700000 - 1600000

= $ 2100000

Now deduct 35% from $2100000

= 2100000 - 735000

= $1365000

Now the state tax is 5% so less state tax too

= 1365000 - 68250

= $1296750

2. State income tax = $ 68750

3. Federal taxable income = $ 2100000

4. Federal income tax = $ 735000

5. Combined tax = 735000 + 68750 = $ 803750

6. Combined tax rate = 803750 ÷ 2100000 × 100 = 38.27 %

7. Combined tax = 2100000 × 38.27% = $803750